### v1

# Universal bound on the Lyapunov spectrum of quantum master equations

Abstract
The spectral properties of positive maps provide pivotal information for understanding the dynamics of quantum systems interacting with their environment. Furthermore, central problems in quantum information such as the characterization of entanglement may be reformulated in terms of spectral properties of positive maps. The present work aims to contribute to a better understanding of the spectrum of positive maps. Specifically, our main result is a new proof of a universal bound on the $d^{2}-1$ generically non vanishing decay rates $\Gamma_{i}$ of time-autonomous quantum master equations on a $d$-dimensional Hilbert space:
$\Gamma_{\mathrm{max}}\,\leq\,c_{d}\,\sum_{i=1}^{d^{2}-1}\Gamma_{i}$
The prefactor $c_{d}$ , which we explicitly determine, depends only on the dimension $d$ and varies depending on the sub-class of positive maps to which the semigroup solution of the master equation belongs. We provide a brief but self-consistent survey of these concepts. We obtain our main result by resorting to the theory of Lyapunov exponents, a central concept in the study of dynamical systems, control theory, and out-of-equilibrium statistical mechanics. We thus show that progress in understanding positive maps in quantum mechanics may require ideas at the crossroads between different disciplines. For this reason, we adopt a notation and presentation style aimed at reaching readers with diverse backgrounds.
1 Introduction
Physics is an experimental science. A recurring question is how to extract dynamical laws from experimental data. In the case of open quantum systems, this question is often formulated by asking whether an observed evolution (dynamical map) can be mathematically understood as a family of linear maps that enjoys the semigroup property and whose members are all positivity preserving. In that case, there also exists a precise and computable answer for time-autonomous dynamics [WoEiCuCi2008]. Specifically, one needs to check whether the logarithm of a tomographically estimated dynamical map at a given time, i.e., a single snapshot of the dynamical map, satisfies the necessary and sufficient conditions characterizing generators of completely positive semigroups and, therefore, solutions of the Lindblad [LinG1976] Gorini-Kossakowski-Sudarshan [GoKoSu1976] master equation. These conditions are preservation of the trace and self-adjointness, as well as conditional complete positivity [EvaD1977] (see also [WolM2012, § 7]) of the generator. Whilst the answer is mathematically complete and constructive, actual implementations encounter several stumbling blocks. To start with, the complex logarithm is a multi-valued function so that, in general, checks may require to extend the search to multiple branches. In addition, the search must be robust against state preparation and measurement errors that occur in realistic situations. Finally, the computational cost grows exponentially with the number of dimensions. Connecting to this cost, it is also worth mentioning that it is in fact possible to prove [CuEiWo2012] that the 11-IN-33-SAT, an instance of a NP-complete problem in computational complexity theory [MoMe2011, § 5], is amenable by a Turing reduction to proving that the exponential of a certain non-representational (i.e., abstract and non related to any physical model [IpCa2021]) generator yields an element of a complete positive semigroup. This means that in principle assessing whether the logarithm of a dynamical map specifies the generator of a completely positive semigroup is at least as hard as solving any NP problem on a classical computer. In practice and for systems with few degrees of freedom, the result of [WoEiCuCi2008] is the basis for computationally viable algorithms which are relevant for to-date quantum computing hardware [LiSeKoOnCu2025].
The difficulties discussed above motivate the interest for partial but practically more direct answers that provide necessary conditions (“witnesses”) to identify a completely positive semigroup. It was conjectured long ago [KimG2002] that complete positivity of a homogeneous semigroup of dynamical maps necessarily entails an upper bound on the relaxation rates characterizing the contractive forward evolution. Relaxation rates are commonly measured in laboratory experiments. The conjecture was later extensively corroborated by proofs under additional hypotheses and examples exhibiting tightness of the putative bound [ChKiKoSh2021]. Mathematically, relaxation rates are the real parts of the logarithms of the eigenvalues of the flow solution of a linear differential equation. If the differential equation is time-autonomous, the forward in time flow it generates constitutes a homogeneous semigroup and the relaxation rates coincide with the Lyapunov exponents [AdrL1995]. These considerations led to a first proof of the conjecture by combining ideas of dynamical systems and open quantum system theory [MGKiCh2025]. The proof, however, left open the question of tight bounds for the relaxation rates of positive but non completely positive maps. This is an important question because positive (i.e., positivity-preserving) linear maps—which are fundamental to quantum mechanics—have a far more articulated structure than those appearing in the theory of classical stochastic processes [ChKo2009]. In [ChvEnKiMG2025, vEnChKiMG2025] we were able to provide a complete answer to this question using different and more traditional methods of algebraic quantum theory. Specifically, the results of [ChvEnKiMG2025, vEnChKiMG2025] show that the same bound holds for maps that are at least $2$-positive and that this requirement cannot be weakened further. We emphasize that $2$-positive maps are completely positive maps for single-qubit systems. The relaxation rates of simply positive (or $1$-positive) maps only obey the trivial bound dictated by contractivity, as do classical positive maps. Finally, our results in [ChvEnKiMG2025, vEnChKiMG2025] single out the class of Schwarz maps, a subset of positive maps containing all $2$-positive maps, by showing that their relaxation rates obey a weaker yet non trivial bound.
The scope of the present contribution is to detail an alternative proof of the results of [ChvEnKiMG2025, vEnChKiMG2025] by revisiting the connection with Lyapunov exponents in dynamical systems theory. We believe this is interesting for several reasons. The advent of nano-technologies has given a lot of impulse to the study of finite-time (stochastic) thermodynamics, thus attracting to the theory of open quantum systems also scholars with multifarious backgrounds and, in particular, statistical physics. The derivation of the results of [ChvEnKiMG2025, vEnChKiMG2025] by directly relating the properties of positive maps in quantum theory with Lyapunov exponents—dynamical systems indicators of cross-disciplinary interest (see, e.g., [CeCeVu2009, PiPo2016, MaShSt2016])—may be useful to a wider community than the one traditionally interested in open quantum systems. In addition, we also show how to use the affine invariance of Lyapunov exponents and the Lozinskii-Dahlquist estimates to derive our main results [LozS1958, DahG1959] (see also [AdrL1995, § IV.6]). These are general results to which the current literature on Lyapunov exponents often pays little attention to (see, however, the recent [AmSo2014, CiVeJaBu2022]), probably as a consequence of the general appeal for computational applications of formulations based on Oseledets’ multiplicative ergodic theorem [OseV1968, BeGaGiSt1980, BeGaGiSt1980a]. Finally, our results provide tight bounds on the maximal Lyapunov exponent of a dynamical system, hence accomplishing a task generically known to be NP-hard [TsBl1997].
With these considerations in mind, we have chosen a presentation style and notation (e.g., linear algebra rather than Dirac’s bra-ket notation from quantum physics) that is likely to be more familiar to readers whose principal research interests lie, for example, in dynamical systems, control theory, and statistical physics.
The structure of the paper is as follows. In section 2 we briefly recall basic facts about positive linear maps in classical and quantum physics. In section 3 we recall that any generator of trace- and self-adjointness preserving linear maps embeds a classical master equation [TeMa1992] if and only if the Kossakowski conditions hold [KosA1972, GoKoSu1976] (see also [RiHu2012, § 4.3]). In section 4 we prove that the foliation of the set of positive maps in quantum mechanics in a hierarchy of subsets translates into conditions on the contraction rate of the volume form of state operators. These results are instrumental to prove the main result of this work. Section 5 contains our main result. We include some examples illustrating the bounds in section LABEL:sec:examples. The last section is devoted to conclusions and outlook.
2 Positive maps solving differential equations
We first recall some facts about classical master equations and how they generate semigroups of positive maps. We then turn to the quantum case to very briefly summarize analogies and differences. Readers familiar with the topics may want to skip this section.
2.1 Classical master equations
Probability vectors valued on a finite-dimensional configuration space are the main objects in the theory of classical stochastic processes. These are real vectors with all positive entries summing up to one. Geometrically they form a simplex, a convex compact subset of a real Euclidean vector space [BeZy2006, § 2].
If the stochastic process is a differentiable Markov chain on a $d$-dimensional Euclidean configuration space, then the family of probability vectors $\mathsf{p}(t)$ that describes the state of the process as $t$ increases, satisfies a first-order system of differential equations of the form
$\dot{\mathsf{p}}(t)=\mathsf{G}(t)\,\mathsf{p}(t)\,.$ (1)
We refer to the family of $d\,\times\,d$-matrices $\mathsf{G}(t)$ as the (time-dependent) generator of the dynamics. We require $\mathsf{G}(t)$ to depend on the time variable $t$ in a sufficiently regular manner to satisfy the conditions of global existence of solutions [AdrL1995, § 1]. In addition, we require $\mathsf{G}(t)$ to be column-diagonally dominated with positive off-diagonal entries, that is,
$\mathsf{G}_{i,j}(t)\begin{cases}\,\geq\,0&\forall\,i\neq j\\[8.5359pt] =-\sum_{k\neq i}^{d}\mathsf{G}_{k,i}(t)&\forall\,i=j\end{cases}$ (2)
Under these conditions, (1) specifies a classical master equation governing the forward in time evolution of probability vectors of the Markov chain. In particular, for each $t\,\geq\,s$ the matrix elements of the time-ordered exponential
$\mathsf{T}(t,s)=\mathcal{T}\exp\left(\int_{s}^{t}\mathrm{d}u\,\mathsf{G}(u)\right)$ (3)
solving (1) with initial condition
$\mathsf{T}(s,s)=\mathsf{1}_{d}$
specify the transition probabilities to go from any state at the earlier time $s$ to another state at the later time $t$. The theory of differential equations guarantees that the maps (3) collectively enjoy the semigroup property
$\mathsf{T}(t_{\mathfrak{3}},t_{\mathfrak{1}})=\mathsf{T}(t_{\mathfrak{3}},t_{\mathfrak{2}})\mathsf{T}(t_{\mathfrak{2}},t_{\mathfrak{1}})\qquad\forall\,t_{\mathfrak{3}}\,\geq\,t_{\mathfrak{2}}\,\geq\,t_{\mathfrak{1}}$
which guarantees that the Chapman-Kolmogorov equation, a necessary consequence of the Markov property, is satisfied [PavG2014, § 2]. This means that the transition probability matrix (3) maps the simplex of probability vectors in $\mathbb{R}^{d}$ into itself. We thus say that (3) defines a family of classical positive linear maps. A further direct consequence of the interpretation of (3) as transition probabilities is that the dynamics generated by (1) is contractive with respect to the $1$-norm
$\|\mathsf{T}(t,s)(\mathsf{p}(s)-\tilde{\mathsf{p}}(s))\|_{\mathfrak{1}}\,\leq\,\left\|\mathsf{p}(s)-\tilde{\mathsf{p}}(s)\right\|_{\mathfrak{1}}\,\equiv\,\sum_{i=1}^{d}|\mathsf{p}_{i}(s)-\tilde{\mathsf{p}}_{i}(s)|\,.$ (4)
Combining contractivity with the property of mapping a convex compact set into itself paves the way for the application of Perron-Frobenius theorem (see e.g [SenE2006, § 1]). The upshot is that the spectrum of any of the semigroup elements (3) is confined to the unit disk of the complex plane, and it is symmetric with respect with the real axis. Furthermore, there is at least one eigenvalue equal to one whose right eigenvector is a probability vector after an appropriate choice of normalization.
2.2 Quantum master equations
State operators are the main objects in the theory of open quantum systems. In the case of a Hilbert space with $d$ dimensions, they are positive semi-definite (henceforth just called “positive”) $d\,\times\,d$ matrices with unit trace. Geometrically they also form a convex compact subspace of the matrix space $\mathcal{M}_{d}(\mathbb{C})$ of $d\,\times\,d$ matrices with complex entries [BeZy2006, § 8.4].
Based only on the requirements of trace and self-adjointness preservation it is in general possible to derive the matrix valued equation
$\dot{\bm{\alpha}}(t)=\mathfrak{L}(t)[\bm{\alpha}(t)]\,.$ (5)
The action of the generator $\mathfrak{L}(t)$ on any $\operatorname{O}\in\mathcal{M}_{d}(\mathbb{C})$ is always amenable to the following canonical form (see, e.g., [HaCrLiAn2014])
$\begin{split}&\mathfrak{L}(t)[\operatorname{O}]=-\imath\,\left[\operatorname{H}(t),\operatorname{O}\right]+\sum_{\ell=1}^{\mathrm{d}^{2}-1}\mathscr{c}_{\ell}(t)\frac{\left[\operatorname{L}_{\ell}(t)\,,\operatorname{O}\operatorname{L}_{\ell}^{\dagger}(t)\right]+\left[\operatorname{L}_{\ell}(t)\operatorname{O}\,,\operatorname{L}_{\ell}^{\dagger}(t)\right]}{2}\\ &\begin{cases}\operatorname{Tr}\left(\operatorname{L}_{\ell}^{\dagger}(t)\operatorname{L}_{k}(t)\right)=\delta_{\ell,k}\\ \operatorname{Tr}\left(\operatorname{L}_{\ell}(t)\right)=0\\ \operatorname{H}(t)=\operatorname{H}^{\dagger}(t)\end{cases}\hskip 14.22636pt\forall\,t\,\geq\,0\hskip 14.22636pt\&\hskip 14.22636pt\forall\,\ell,k=1,\dots,d^{2}-1\end{split}$ (6)
where the $\mathscr{c}_{\ell}(t)$’s are scalar couplings and the $\operatorname{L}_{\ell}(t)$ are commonly referred to as decoherence or jump operators. As in the classical case, we assume that any explicit time dependence is sufficiently regular to satisfy the conditions for global existence of solutions of (5). In particular, here we require
$|c_{\ell}(t)|\,<\,\infty\qquad\forall\,\ell=1,\dots,d^{2}-1$
We refer to (5), always implying the generator form (6), as “quantum master equation”. The main problem with (5) is that it does not ensure positivity preservation without further assumptions on the scalar couplings $\mathscr{c}_{\ell}(t)$.
A cornerstone result of the theory of open quantum systems is the proof first obtained in the time-autonomous case by Lindblad [LinG1976] and Gorini-Kossakowski-Sudarshan [GoKoSu1976], that if and only if
$\mathscr{c}_{\ell}(t)\,\geq\,0\qquad\forall\,\ell=1,\dots,d^{2}$ (7)
then solutions of (5) can always be couched into the readily positivity-preserving operator sum form (often referred to as Kraus form [KraK1983])
$\Phi(t,s)[\bm{\alpha}(s)]=\sum_{\ell=0}^{d^{2}-1}\operatorname{V}_{\ell}(t,s)\bm{\alpha}(s)\operatorname{V}_{\ell}^{\dagger}(t,s)\qquad\forall\,t\,\geq\,s$ (8)
The collection of matrices $\operatorname{V}_{\ell}(t,s)\in\mathcal{M}_{d}(\mathbb{C})$ is such that (8) reduces to the identity map for $t=s$ and
$\sum_{\ell=0}^{d^{2}-1}\operatorname{V}_{\ell}^{\dagger}(t,s)\operatorname{V}_{\ell}(t,s)=\operatorname{1}_{d}\qquad\forall\,t\,\geq\,s$
to ensure trace preservation.
It is worth recalling that we can always couch the matrix equation (5) into the form of a vector equation using the vectorization isomorphism relating outer and tensor products [BeZy2006, § 8.2]. Accordingly, a solution of (5) is in one-to-one correspondence with a family of vectors in $\mathbb{C}^{d^{2}}$
$\bm{\alpha}(t)=\sum_{i,j=1}^{d}\mathbf{e}_{i}^{\scriptscriptstyle{(d)}}\mathbf{e}_{j}^{\scriptscriptstyle{(d)}}{}^{\dagger}\alpha_{i,j}(t)\quad\Longleftrightarrow\quad\mathscr{a}(t)=\sum_{i,j=1}^{d}\mathbf{e}_{i}^{\scriptscriptstyle{(d)}}\otimes\mathbf{e}_{j}^{\scriptscriptstyle{(d)}}\alpha_{i,j}(t)$ (9)
where $\left\{\mathbf{e}_{i}^{\scriptscriptstyle{(d)}}\right\}_{i=1}^{d}$ is the canonical (or computational) basis of $\mathbb{C}^{d}$. In words, (9) says that we can construct a vector out of a matrix by stacking rows one above the other. Upon applying this prescription and using the overline symbol $\overline{(\cdot)}\,$ to denote entry-wise complex conjugation, the evolution map (8) takes the familiar form in linear dynamical systems theory (see, e.g., [CeCeVu2009]) of a matrix acting on a vector:
$\mathscr{F}(t,s)\mathscr{a}(s)=\left(\sum_{\ell=0}^{d^{2}-1}\operatorname{V}_{\ell}(t,s)\otimes\overline{\operatorname{V}_{\ell}}(t,s)\right)\mathscr{a}(s)\qquad\forall\,t\,\geq\,s$
The $d^{2}\,\times\,d^{2}$ matrices $\mathscr{F}(t,s)$ are the elements of the semigroup solving the system of differential equations
$\begin{split}&\partial_{t}\mathscr{F}(t,s)=\mathscr{L}(t)\mathscr{F}(t,s)\\ &\mathscr{F}(s,s)=\mathsf{1}_{d^{2}}\end{split}\qquad\forall\,t\,\geq\,s$ (10)
with generator
$\mathscr{L}(t)$ $=-\imath\,\Big(\operatorname{H}(t)\otimes\mathsf{1}_{d}-\mathsf{1}_{d}\otimes\operatorname{H}^{\top}(t)\Big)$
$+\sum_{\ell=1}^{\mathrm{d}^{2}-1}\mathscr{c}_{\ell}(t)\left(\operatorname{L}_{\ell}(t)\otimes\overline{\operatorname{L}_{\ell}}(t)-\frac{\operatorname{L}_{\ell}^{\dagger}(t)\operatorname{L}_{\ell}(t)\otimes\mathsf{1}_{d}+\mathsf{1}_{d}\otimes\operatorname{L}_{\ell}^{\top}(t)\overline{\operatorname{L}_{\ell}}(t)}{2}\right)$
where $\top$ denotes matrix transposition. The one-to-one correspondence between (10) and (7) implies that also the operator sum (8) enjoys the semigroup property:
$\Phi(t_{\mathfrak{3}},t_{\mathfrak{2}})\big[\Phi(t_{\mathfrak{2}},t_{\mathfrak{1}})[\operatorname{O}]\big]=\Phi(t_{\mathfrak{3}},t_{\mathfrak{1}})[\operatorname{O}]\qquad\forall\,t_{\mathfrak{3}}\,\geq\,t_{\mathfrak{2}}\,\geq\,t_{\mathfrak{1}}\quad\&\quad\forall\,\operatorname{O}\in\mathcal{M}_{d}(\mathbb{C})$
The semigroup generated by the quantum master equation (5) subject to the conditions (7) always preserves state operators. All these facts combined are reminiscent of the properties of semigroups generated by the classical master equation (1). For this reason quantum master equations satisfying (7) often called “Markovian”. The set of positive maps in quantum theory is, however, far more involved than in the classical case. The positive operator sum form (8) is only a sufficient condition for positivity preservation. In particular, it is equivalent to adding to positivity preservation the requirement [StiW1955, ChoM1975] that the map on $\mathcal{M}_{d^{2}}(\mathbb{C})$ obtained by the tensor product of (8) with the identity on $\mathcal{M}_{d}(\mathbb{C})$ is also positive:
$\tilde{\Phi}^{(d)}(t,s)[\tilde{\operatorname{O}}]:=(\mathrm{Id}_{d}\otimes\Phi(t,s))[\tilde{\operatorname{O}}]\,\geq\,0\qquad\forall\,t\,\geq\,s$ (11)
for all positive matrices $\tilde{\operatorname{O}}$ in $\mathcal{M}_{d^{2}}(\mathbb{C})$. The condition (11) is called $d$- or complete positivity. As a consequence, the master equation (1) subject to the Lindblad-Gorini-Kossakowski-Sudarshan conditions (7) may be more descriptively called completely positive rather than Markovian in reference to the property of the solution semigroup [BaNaTh2008].
Inspecting (11) immediately suggests the existence of a hierarchy of $k$-positive maps $\tilde{\Phi}^{(k)}(t,s)$, $k=1,\dots,d$ characterized by the conditions
$\tilde{\Phi}^{(k)}(t,s)[\tilde{\operatorname{O}}]:=(\mathrm{Id}_{k}\otimes\Phi(t,s))[\tilde{\operatorname{O}}]\,\geq\,0\qquad\forall\,t\,\geq\,s$
where now $\mathrm{Id}_{k}$ is the identity on $\mathcal{M}_{k}(\mathbb{C})$ and the inequality holds for all positive matrices $\tilde{\operatorname{O}}$ in $\mathcal{M}_{kd}(\mathbb{C})$. If we then denote by $\mathcal{P}_{k}$ the set of $k$-positive maps, it is rather straightforward to prove the chain of inclusions [ChrD2022]
$\mathcal{P}\,\equiv\,\mathcal{P}_{\mathfrak{1}}\supset\mathcal{P}_{\mathfrak{2}}\supset\dots\supset\mathcal{P}_{d}\,\equiv\,\mathcal{CP}$ (12)
We thus see that the class of $d$-positive maps, i.e., of completely positive maps ($\mathcal{CP}$), is the smallest set in the above chain of inclusions. Unfortunately, for $k$-positive maps there is no general counterpart of the operator-sum representation (8) that characterizes completely positive maps. Analogously, on the differential level the conditions (7) have no general counterpart to specify generators of $k$-positive semigroups [WoCi2008, ChrD2022]. Even if it can be argued that completely positive maps are the only ones adapted to describe the evolution of open quantum systems (see, however, the debate [PecP1994, AliR1995, PecP1995] and the more recent [DoLi2016]), completely positive dynamical maps can be solutions of a quantum master equation (5) whose generator does not satisfy the complete positivity conditions (7). The upshot is that quantum master equations not satisfying (7) must be necessarily considered. We refer to [SmVa2010] for an example with clear physical interpretation. In addition, understanding the properties of $k$-positive semigroups for all $k$ is necessary to determine criteria to discriminate between entangled and separable states (see [ChKo2009] and references therein).
We conclude the section recalling that a trace and self-adjointness preserving linear map $\Phi$ is positive if and only if is a contraction [KosA1972]
$\left\|\Phi(\operatorname{O})\right\|_{\mathfrak{1}}\,\leq\,\left\|\operatorname{O}\right\|_{\mathfrak{1}}\qquad\forall\,\operatorname{O}=\operatorname{O}^{\dagger}\in\mathcal{M}_{d}(\mathbb{C})$
with respect the trace norm (also known as Schatten $1$-norm)
$\left\|\operatorname{O}\right\|_{\mathfrak{1}}:=\operatorname{Tr}\left(\sqrt{\operatorname{O}^{\dagger}\operatorname{O}}\right)\,.$
The trace norm for self-adjoint operators coincides with the sum over the absolute value of all eigenvalues. Thus, positive linear maps are contractive and transform a convex compact set into itself. Again, these facts permit to demonstrate that positive maps satisfy the hypotheses of the Perron-Frobenius theorem. The conclusion is that all the eigenvalues belong to the unit disk of the complex plane and are symmetric with respect with the real axis. Furthermore, there is at least one eigenvalue equal to one whose right eigenoperator is a state operator under suitable normalization.
3 Embedded classical master equation
As (5) is self-adjointness preserving, solutions that evolve from self-adjoint initial conditions at $t=0$
$\bm{\alpha}(0)=\sum_{i=1}^{d}p_{i}\,\bm{\varphi}_{i}\bm{\varphi}_{i}^{\dagger}$ (13)
at any later instant of time are amenable to the diagonal form
$\bm{\alpha}(t)=\sum_{i=1}^{d}\wp_{i}(t)\,\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)$ (14)
where both the $\bm{\varphi}_{i}$’s and the $\bm{\phi}_{\mathfrak{i}}(t)$’s constitute a complete orthonormal basis of $\mathbb{C}^{d}$ and the $p_{i}$’s and $\wp_{i}$’s respectively denote the eigenvalues of $\bm{\alpha}(0)$, $\bm{\alpha}(t)$.
As orthonormal bases are connected by unitary maps, there exists a family of unitaries such that
$\bm{\phi}_{i}(t)=\mathsf{U}(t)\bm{\varphi}_{i}$ (15)
If we insert (15) into (5) and denote by $\bm{p}$, $\bm{\wp}(t)$ the vectors whose entries are the eigenvalues of (13), (14) respectively, a straightforward computation [TeMa1992, WiTo1999] yields
$\begin{split}&\bm{\dot{\wp}}(t)=\mathsf{W}(t)\,\bm{\wp}(t)\\ &\bm{\wp}(0)=\bm{p}\,.\end{split}$ (16)
At any time $t$ the $d\times d$ matrices $\mathsf{W}(t)$ have the same structure as the generator of the classical master equation (2): they are now column-diagonally dominated with elements
$\mathsf{W}_{i,j}(t)=\left\langle\,\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,,\mathfrak{L}(t)[\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)]\,\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,\right\rangle=\begin{cases}\mathsf{R}_{i,j}(t)&\forall\,i\neq j\\[8.5359pt] -\sum_{k\neq i}^{d}\mathsf{R}_{i,j}(t)&\forall\,i=j\end{cases}$ (17)
where $\left\langle\,\cdot\,$$\cdot\,\right\rangle$ is the standard inner product in $\mathbb{C}^{d}$ and the off-diagonal elements are the sum over the corresponding matrix elements of the decoherence operators weighted by the canonical couplings:
$\mathsf{R}_{i,j}(t)=\sum_{\ell=1}^{d^{2}-1}\mathscr{c}_{\ell}(t)\,\left|\left\langle\,\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,,\operatorname{L}_{\ell}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\,\right\rangle\right|^{2}$ (18)
Before we continue some remarks are in order.
  • In general the “spectral flow” equations (15), (16) do not provide a constructive method to solve (5) but are only a tool to analyze the kinematic properties of individual solutions. More precisely, the matrix elements of (18) depend upon the unitary evolution operator in (15). The only general method to identify this latter one requires having first solved (5) for a given self-adjoint initial condition (13). This also means that the family of matrices $\mathsf{U}(t)$ and $\mathsf{W}(t)$ depend upon the choice of initial conditions. In other words, the spectral flow equations play for the kinematic analysis of quantum master equations the same role as the current velocity formalism for the analysis of the Fokker-Planck equation [NelE2001, § 13].
  • The derivation of (16) requires almost-everywhere differentiability of eigenvalues and eigenvectors of families of self-adjoint operators satisfying (5). Mathematically, any solution $\mathscr{x}(t)$ of a linear differential equation with respect to a real variable $t$ can be analytically continued to a unique holomorphic function of the independent variable:
    $\overline{\mathscr{x}(z)}=\mathscr{x}(\overline{z})\qquad\forall\,z\in\mathbb{C}\quad\text{ where }\operatorname{Re}z=t\,.$
    As a consequence, we can analytically continue self-adjoint solutions of (5) to matrix valued holomorphic functions
    $\bm{\alpha}^{\dagger}(z)=\bm{\alpha}(\bar{z})$ (19)
    in a neighborhood of the real $t$-axis. This fact allows us to invoke Theorem 6.1 in [KatT1982, § 2.6.1] guaranteeing that eigenvalues and spectral projectors of a family of operators enjoying (19) are also holomorphic functions and as such are continuous and differentiable. In addition [KatT1982, § 2.6.1], eigenvectors can also be chosen to be holomorphic functions, thus ensuring the validity of the construction used in the proof.
3.1 Kossakowski conditions
The representation (14) yields the identity
$\left\|\bm{\alpha}(t)\right\|_{\mathfrak{1}}=\sum_{i=1}^{d}|\wp_{i}(t)|=\left\|\bm{\wp}(t)\right\|_{\mathfrak{1}}$ (20)
holding for any solution of (5) with self-adjoint initial data. This means that (5) is contractive if and only if all classical master equations embedded in particular solutions are contractive.
It is readily clear that if the complete positivity conditions (7) holds true then the off-diagonal entries (18) are positive functions of time thus implying that (16) is a classical master equation, and, as such describes contractive dynamics. Conversely, if the (18) are positive functions of time for any solution of (5) corresponding to self-adjoint initial data, then the semigroup generated by (5) is necessarily contractive and therefore positive [ChrD2022]. In fact, requiring that the (18) are always positive is equivalent to the Kossakowski conditions [KosA1972, GoKoSu1976] (see also [RiHu2012, § 4.3]) asserting that a linear map of the form (6) is the generator of a trace-preserving contraction semigroup if and only if its matrix entries evaluated on an arbitrary orthonormal frame specify a classical generator. More explicitly, if we denote by $\big\{\operatorname{P}_{i}\big\}_{i=1}^{d}$ a complete set of orthonormal projectors on $\mathbb{C}^{d}$ (that is,
$\sum_{i=1}^{d}\operatorname{P}_{i}=\operatorname{1}_{d}\&$ $\operatorname{P}_{i}\operatorname{P}_{j}=\delta_{i,j}\operatorname{P}_{i}\qquad\forall\,i,j=1,\dots,d\,)$
then the Kossakowski conditions read
$\begin{cases}\operatorname{Tr}\left(\operatorname{P}_{i}\mathfrak{L}(t)[\operatorname{P}_{j}]\right)\,\geq\,0&i\neq j\\[8.5359pt] \operatorname{Tr}\left(\operatorname{P}_{i}\mathfrak{L}(t)[\operatorname{P}_{i}]\right)=-\sum_{k\neq i=1}^{d}\operatorname{Tr}\left(\operatorname{P}_{k}\mathfrak{L}(t)[\operatorname{P}_{i}]\right)\,\leq\,0&i=j\end{cases}$
To summarize: the representation of solutions of the matrix equation (5) in terms of the eigenvector (15) and eigenvalue flow (15) equations provide us with implicit conditions to assess whether (5) generates a positive semigroup.
4 Conditions on the embedded classical master equation stemming from the positivity class of the semigroup
We now turn to derive some properties of (16) that follow if we assume that the semigroup, in addition to being positive and trace-preserving, enjoys further properties. To this end, we recall that for any a complete orthonormal basis $\big\{\bm{f}_{i}\big\}_{i=1}^{d}$ of $\mathbb{C}^{d}$ the generator of a positive semigroup also satisfies
$\operatorname{Tr}\mathfrak{L}(t):=\sum_{i,j=1}^{d}\left\langle\,\bm{f}_{i}\,,\mathfrak{L}(t)[\bm{f}_{i}\bm{f}_{i}^{\dagger}]\,\bm{f}_{j}\,\right\rangle=\operatorname{Tr}\mathscr{L}(t)=-d\,\sum_{\ell=1}^{d^{2}-1}\mathscr{c}_{\ell}(t)\,\leq\,0$ (21)
which implies contractivity of the volume element for any $t\,\geq\,s$:
$\det\mathscr{F}(t,s)=\exp\left(\int_{s}^{t}\mathrm{d}u\,\operatorname{Tr}\mathscr{L}(u)\right)$ (22)
where $\mathscr{F}(t,s)$ is the solution of (10).
4.1 Condition on the embedded classical master equation imposed by $2$-positivity
A linear map
$\bm{X}\colon\mathcal{M}_{d}(\mathbb{C})\mapsto\mathcal{M}_{d}(\mathbb{C})$
is called $2$-positive if for any positive operator $\operatorname{O}$ belonging to $\mathcal{M}_{2d}(\mathbb{C})$
$\operatorname{O}\,\geq\,0\quad\Longrightarrow\quad(\mathrm{Id}_{2}\otimes\bm{X})(\operatorname{O})\,\geq\,0\,.$
Let us now assume that (5) generates $2$-positive semigroup. Then we claim that the classical master equation constructed out of any of the particular solutions is such that
$0\,\geq\,\operatorname{Tr}\mathsf{W}(t)\,\geq\,\frac{1}{d}\operatorname{Tr}\mathfrak{L}(t)$ (23)
An equivalent version of this claim was first proven in [vEnChKiMG2025].
4.1.1 Proof of the claim
The action of a semigroup element close to the identity on a state operator $\bm{\rho}$ admits the explicit representation
$\Phi(t+\varepsilon,t)[\bm{\rho}]=\bm{\rho}+\varepsilon\,\mathfrak{L}(t)[\bm{\rho}]+O(\varepsilon^{2})$
If we pick a vector in $\mathbb{C}^{2d}$ of the form
$\bm{u}=\frac{\mathbf{e}_{1}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}^{\scriptscriptstyle{(i)}}(t)+\mathbf{e}_{2}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}^{\scriptscriptstyle{(j)}}(t)}{\sqrt{2}}$
where $\left\{\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\right\}_{i=1}^{d}$ is a complete orthonormal basis of $\mathbb{C}^{d}$ determined by the eigenvector flow equation (15). The projector along $\bm{u}$ is clearly a positive matrix
$\operatorname{P}_{u}=\bm{u}\bm{u}^{\dagger}=\frac{1}{2}\begin{bmatrix}\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)&\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\\ \bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)&\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\end{bmatrix}$.
Then we find
$(\mathrm{Id}_{2}\otimes\Phi(t+\varepsilon,t))[\operatorname{P}_{u}]$ $=\frac{1}{2}\begin{bmatrix}\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)&\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\\ \bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)&\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\end{bmatrix}$
$+\frac{\varepsilon}{2}\begin{bmatrix}\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)\big]&\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\big]\\ \mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)\big]&\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\big]\end{bmatrix}+O(\varepsilon^{2})\,.$
As the semigroup is $2$-positive by assumption, we must have
$\left\langle\,\bm{v}\,,(\mathrm{Id}_{2}\otimes\Phi(t+\varepsilon,t))[\operatorname{P}_{u}]\,\bm{v}\,\right\rangle\,\geq\,0$ $\forall\,\bm{v}\in\,\mathbb{C}^{2d}$
In particular, we can choose
$\bm{v}=\frac{\mathbf{e}_{1}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}^{\scriptscriptstyle{(i)}}(t)-\mathbf{e}_{2}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}^{\scriptscriptstyle{(j)}}(t)}{\sqrt{2}}=\frac{1}{\sqrt{2}}\begin{bmatrix}\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\\ -\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\end{bmatrix}$
so that
$\left\langle\,\bm{u}\,,\bm{v}\,\right\rangle=0\,.$
In such a case, after a straightforward calculation we obtain
$0\,\leq\,\lim_{\varepsilon\searrow 0}\frac{\left\langle\,\bm{v}\,,(\mathrm{Id}_{2}\otimes\Phi(t+\varepsilon,t))[\operatorname{P}_{u}]\bm{v}\,\right\rangle}{\varepsilon}$
$\hskip 14.22636pt=\frac{\left\langle\,\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)\big]\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,\right\rangle+\left\langle\,\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)\big]\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\,\right\rangle}{2}$
$\hskip 14.22636pt-\frac{\left\langle\,\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\big]\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\,\right\rangle+\left\langle\,\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)\big]\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,\right\rangle}{2}$
Summing over $i,j=1,\dots,d$, the definitions (17) and (21) yield the claim.
4.2 Condition on the embedded classical master equation imposed by the Schwarz-class
It can be shown (see, e.g., the discussion in [ChKiMu2024] and references therein) that the convex set of positive maps on $\mathcal{M}_{d}(\mathbb{C})$ yields a hierarchy of inclusions finer than (12). For the present work we only notice that between the convex sets of positive and $2$-positive maps sits the set $\mathcal{S}_{1}$ (also convex) of Schwarz maps:
$\mathcal{P}\,\equiv\,\mathcal{P}_{\mathfrak{1}}\supset\mathcal{S}_{1}\supset\mathcal{P}_{\mathfrak{2}}$
We characterize Schwarz maps by considering their adjoint with respect to the Hilbert-Schmidt inner product. Specifically, if $\bm{X}$ is an arbitrary linear map, we define its Hilbert-Schmidt adjoint $\bm{X}^{\ddagger}$ by requiring
$\operatorname{Tr}\left(\operatorname{A}^{\dagger}\bm{X}[\operatorname{B}]\right)=\operatorname{Tr}\left(\big(\bm{X}^{\ddagger}[\operatorname{A}]\big)^{\dagger}\operatorname{B}\right)\qquad\forall\,\operatorname{A},\operatorname{B}\in\mathcal{M}_{d}(\mathbb{C})$
It is straightforward to verify that trace-preservation imposes the adjoint to be unital
$\bm{X}^{\ddagger}[\operatorname{1}_{d}]=\operatorname{1}_{d}\,.$
We say that a positive unital map is Schwarz class if
$\bm{X}^{\ddagger}[\operatorname{A}^{\dagger}\operatorname{A}]\,\geq\,\bm{X}^{\ddagger}[\operatorname{A}^{\dagger}]\bm{X}^{\ddagger}[\operatorname{A}]\qquad\forall\,\operatorname{A}\,\in\,\mathcal{M}_{d}(\mathbb{C})$
It is worth recalling here that Kadison’s theorem asserts that the same inequality is satisfied by all positive maps if the domain is restricted to normal operators [KadR1952]. Conversely, $2$-positivity of unital maps implies the Schwarz property.
Let us assume that a semigroup is Schwarz class. A straightforward calculation shows that the generator of such a semigroup must satisfy the dissipativity condition:
$\mathfrak{L}^{\ddagger}(t)[\operatorname{A}^{\dagger}\operatorname{A}]\,\geq\,\mathfrak{L}^{\ddagger}(t)[\operatorname{A}^{\dagger}]\operatorname{A}+\operatorname{A}^{\dagger}\mathfrak{L}^{\ddagger}(t)[\operatorname{A}]\qquad\forall\,t\quad\&\quad\forall\,\,\operatorname{A}\,\in\,\mathcal{M}_{d}(\mathbb{C})\,.$ (24)
We refer to [ChKiMu2024] for explicit examples.
Let us now suppose that that (5) generates a semigroup in $\mathcal{S}_{1}$. Then we claim that the classical master equation constructed out of any of the particular solutions is such that
$0\,\geq\,\operatorname{Tr}\mathsf{W}(t)\,\geq\,\frac{2}{d+1}\operatorname{Tr}\mathscr{L}(t)\,.$ (25)
An equivalent version of this claim was first proven in [ChvEnKiMG2025].
4.2.1 Proof of the claim
If we evaluate the dissipativity condition (24) on non-self-adjoint projectors constructed with eigenvectors of (14) at time $t$
$\operatorname{P}_{i,j}(t)=\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)$
we get
$\mathfrak{L}(t)^{\ddagger}[\operatorname{P}_{j,j}(t)]\,\geq\,\operatorname{P}_{i,j}^{\dagger}(t)\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}(t)]+\mathfrak{L}(t)^{\ddagger}[\operatorname{P}_{i,j}^{\dagger}(t)]\operatorname{P}_{i,j}(t)\,.$
From the left-hand side of the inequality we get
$\sum_{i=1}^{d}\sum_{j\neq i}^{d}\operatorname{Tr}\left(\operatorname{P}_{j,j}(t)\mathfrak{L}(t)^{\ddagger}\big[\operatorname{P}_{j,j}(t)\big]\right)=\sum_{i,j=1}^{d}(1-\delta_{i,j})\operatorname{Tr}\left(\mathfrak{L}(t)\big[\operatorname{P}_{j,j}(t)\big]\operatorname{P}_{j,j}(t)\right)=(d-1)\,\operatorname{Tr}\mathsf{W}(t)$
The result must be larger than or equal to the corresponding expression for the right-hand side of (24)
$\sum_{i=1}^{d}\sum_{j\neq i}^{d}\operatorname{Tr}\left(\operatorname{P}_{j,j}(t)\big(\operatorname{P}_{i,j}^{\dagger}(t)\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}(t)]+\mathfrak{L}(t)^{\ddagger}[\operatorname{P}_{i,j}^{\dagger}(t)]\operatorname{P}_{i,j}(t)\big)\right)$
$=\sum_{i,j=1}^{d}(1-\delta_{i,j})\operatorname{Tr}\left(\operatorname{P}_{j,i}(t)\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}(t)]+\mathfrak{L}(t)^{\ddagger}[\operatorname{P}_{i,j}^{\dagger}(t)]\operatorname{P}_{i,j}(t)\right)=2\,\left(\operatorname{Tr}\mathscr{L}(t)-\operatorname{Tr}\mathsf{W}(t)\right)$
The claim follows at once by comparing the two above identities.
5 Main result
The Lyapunov characteristic exponents $\big\{\lambda_{i}\big\}_{i=0}^{d^{2}-1}$ (see, e.g., [OseV1968, BeGaGiSt1980] or the monographs [CeCeVu2009, PiPo2016]; a brief summary is also in  LABEL:app:Lyapunov below) of a quantum master equation specify the asymptotic properties of averages over a state operator. As such they are also experimentally accessible. The Lyapunov exponents of a trace- and self-adjointness preserving, contractive (also referred to as positive) semigroup solving (5) satisfy
$0\,\geq\,\lambda_{\mathfrak{0}}\,\geq\,\lambda_{\mathfrak{1}}\,\geq\,\dots\,\geq\,\lambda_{d^{2}-1}$
We are now in a position to prove results equivalent to those of [ChvEnKiMG2025, vEnChKiMG2025].
Proposition 5.1
If a semigroup generated by (5) is trace-preserving, self-adjointness preserving, contractive,
  1. 1.
    and also 22-positive, i.e., (23) holds, then the smallest Lyapunov exponent obeys the bound
    $\lambda_{d^{2}-1}\,\geq\,\frac{1}{d}\sum_{i=1}^{d^{2}-1}\lambda_{i}$ (26)
  2. 2.
    and of Schwarz class, i.e., (25) holds, then
    $\lambda_{d^{2}-1}\,\geq\,\frac{2}{d+1}\sum_{i=1}^{d^{2}-1}\lambda_{i}$ (27)
  3. 3.
    but does not enjoy any additional property, then the lower bound on the smallest Lyapunov exponent is trivial:
    $\lambda_{d^{2}-1}\,\geq\,\sum_{i=1}^{d^{2}-1}\lambda_{i}$ (28)
Before turning to the proof some observations are in order.
  • The set of Lyapunov exponents identifies the asymptotic decay rates of a contractive semigroup via
    $\Gamma_{\mathfrak{i}}=-\lambda_{\mathfrak{i}}$ $\mathfrak{i}=0,\dots,d^{2}-1$
    In terms of the asymptotic decay rates the inequalities read
    $\Gamma_{d^{2}-1}\leq c_{d}\sum_{\ell=1}^{d^{2}-1}\Gamma_{\ell}$
    where the universal prefactor $c_{d}$ is
    $c_{d}=\begin{cases}\end{cases}$