### v2

# Universal bound on the Lyapunov spectrum of quantum master equations

Abstract
The spectral properties of positive maps are pivotal for understanding the dynamics of quantum systems interacting with their environment. Furthermore, central problems in quantum information such as the characterization of entanglement can be reformulated in terms of spectral properties of positive maps. The present work aims to contribute to a better understanding of the spectrum of positive maps. Specifically, our main result is a new proof of a universal bound on the $d^{2}-1$ generically non vanishing decay rates $\Gamma_{i}$ of time-autonomous quantum master equations on a $d$-dimensional Hilbert space:
$\phantom{aaaaaaaaaaaaaa}\Gamma_{\mathrm{max}}\,\leq\,\varkappa_{d}\,\sum_{i=1}^{d^{2}-1}\Gamma_{i}$
The prefactor $\varkappa_{d}$ depends only on the dimension $d$ and varies depending on the sub-class of positive maps to which the semigroup solution of the master equation belongs. We provide a brief but self-consistent survey of these concepts. We obtain our main result by resorting to the theory of Lyapunov exponents, a central concept in the study of dynamical systems, control theory, and out-of-equilibrium statistical mechanics. We thus show that progress in understanding positive maps in quantum mechanics may require ideas at the crossroads between different disciplines. For this reason, we adopt a notation and presentation style aimed at reaching readers with diverse backgrounds.
1 Introduction
Physics is an experimental science. A recurring question is how to extract dynamical laws from experimental data. In the case of open quantum systems, this question is often formulated by asking whether an observed evolution (dynamical map) can be mathematically understood as a family of linear maps that enjoys the semigroup property and whose members are all positivity preserving. In that case, there also exists a precise and computable answer for time-autonomous dynamics [WoEiCuCi2008]. Specifically, one needs to check whether the logarithm of a tomographically estimated dynamical map at a given time, i.e., a single snapshot of the dynamical map, satisfies the necessary and sufficient conditions characterizing generators of completely positive semigroups and, therefore, solutions of the Lindblad [LinG1976] Gorini-Kossakowski-Sudarshan [GoKoSu1976] master equation. These conditions are preservation of the trace and self-adjointness, as well as conditional complete positivity [EvaD1977] (see also [WolM2012, § 7]) of the generator. Whilst the answer is mathematically complete and constructive, actual implementations encounter several stumbling blocks. To start with, the complex logarithm is a multi-valued function so that, in general, checks may require to extend the search to multiple branches. In addition, the search must be robust against state preparation and measurement errors that occur in realistic situations. Finally, the computational cost grows exponentially with the number of dimensions. Connecting to this cost, it is also worth mentioning that it is in fact possible to prove [CuEiWo2012] that the 11-IN-33-SAT, an instance of a NP-complete problem in computational complexity theory [MoMe2011, § 5], is amenable by a Turing reduction to proving that the exponential of a certain non-representational (i.e., abstract and non related to any physical model [IpCa2021]) generator yields an element of a completely positive semigroup. This means that in principle assessing whether the logarithm of a dynamical map specifies the generator of a completely positive semigroup is at least as hard as solving any NP problem on a classical computer. In practice and for systems with few degrees of freedom, the result of [WoEiCuCi2008] is the basis for computationally viable algorithms which are relevant for to-date quantum computing hardware [LiSeKoOnCu2025].
The difficulties discussed above motivate the interest for partial but practically more direct answers that provide necessary conditions (“witnesses”) to identify a completely positive semigroup. It was conjectured long ago [KimG2002] that complete positivity of a homogeneous semigroup of dynamical maps necessarily entails an upper bound on the relaxation rates characterizing the contractive forward evolution. Relaxation rates are commonly measured in laboratory experiments. The conjecture was later extensively corroborated by proofs under additional hypotheses and examples exhibiting tightness of the putative bound [ChKiKoSh2021]. Mathematically, relaxation rates are the real parts of the logarithms of the eigenvalues of the flow solution of a linear differential equation. If the differential equation is time-autonomous, the forward in time flow it generates constitutes a homogeneous semigroup and the relaxation rates coincide with the Lyapunov exponents [AdrL1995]. These considerations led to a first proof of the conjecture by combining ideas of dynamical systems and open quantum system theory [MGKiCh2025]. The proof, however, left open the question of tight bounds for the relaxation rates of positive but non completely positive maps. This is an important question because positive (i.e., positivity-preserving) linear maps—which are fundamental to quantum mechanics—have a far more articulated structure than those appearing in the theory of classical stochastic processes [ChKo2009]. In [ChvEnKiMG2025, vEnChKiMG2025] we were able to provide a complete answer to this question using different and more traditional methods of algebraic quantum theory. Specifically, the results of [ChvEnKiMG2025, vEnChKiMG2025] show that the same bound holds for maps that are at least $2$-positive and that this requirement cannot be weakened further. We emphasize that $2$-positive maps are completely positive maps for single-qubit systems. The relaxation rates of simply positive (or $1$-positive) maps only obey the trivial bound dictated by contractivity, as do classical positive maps. Finally, our results in [ChvEnKiMG2025, vEnChKiMG2025] single out the class of Schwarz maps, a subset of positive maps that contains the class of $2$-positive maps: we show that the relaxation rates of Schwarz maps obey a weaker yet non trivial bound than that of $2$-positive maps.
The scope of the present contribution is to detail an alternative proof of the results of [ChvEnKiMG2025, vEnChKiMG2025] by revisiting the connection with Lyapunov exponents in dynamical systems theory. We believe this is interesting for several reasons. The advent of nano-technologies has given a lot of impulse to the study of finite-time (stochastic) thermodynamics, thus attracting to the theory of open quantum systems also scholars with multifarious backgrounds and, in particular, statistical physics. The derivation of the results of [ChvEnKiMG2025, vEnChKiMG2025] by directly relating the properties of positive maps in quantum theory with Lyapunov exponents—dynamical systems indicators of cross-disciplinary interest (see, e.g., [CeCeVu2009, PiPo2016, MaShSt2016])—may be useful to a wider community than the one traditionally interested in open quantum systems. In addition, we also show how to use the affine invariance of Lyapunov exponents and the Lozinskii-Dahlquist estimates to derive our main results [LozS1958, DahG1959] (see also [AdrL1995, § IV.6]). These are general results to which the current literature on Lyapunov exponents often pays little attention to (see, however, the recent [AmSo2014, CiVeJaBu2022]), probably as a consequence of the general appeal for computational applications of formulations based on Oseledets’ multiplicative ergodic theorem [OseV1968, BeGaGiSt1980, BeGaGiSt1980a]. Finally, our results provide tight bounds on the maximal Lyapunov exponent of a dynamical system, hence accomplishing a task generically known to be NP-hard [TsBl1997].
With these considerations in mind, we have chosen a presentation style and notation (e.g., linear algebra rather than Dirac’s bra-ket notation from quantum physics) that is likely to be more familiar to readers whose principal research interests lie, for example, in dynamical systems, control theory, and statistical physics.
The structure of the paper is as follows. In section 2 we briefly recall basic facts about positive linear maps in classical and quantum physics. In section 3 we recall that any generator of trace- and self-adjointness preserving linear maps embeds a classical master equation [TeMa1992] if and only if the Kossakowski conditions hold [KosA1972, GoKoSu1976] (see also [RiHu2012, § 4.3]). In section 4 we prove that the foliation of the set of positive maps in quantum mechanics in a hierarchy of nested subsets translates into conditions on the contraction rate of the volume form of state operators. These results are instrumental to prove the main result of this work. In section 5 we briefly review basic concepts of the theory of Lyapunov exponents that we use in the proof of our main result. We give particular emphasis to Lyapunov regularity, a property that quantum master equations relevant for applications, most often time-autonomous or time-(almost)-periodic, enjoy. Section 6 contains our main result. We include some examples illustrating the bounds in section LABEL:sec:examples. The last section is devoted to conclusions and outlook.
2 Positive maps solving differential equations
We first recall some facts about classical master equations and how they generate semigroups of positive maps. We then turn to the quantum case to very briefly summarize analogies and differences. Readers familiar with the topics may want to skip this section.
2.1 Classical master equations
Probability vectors valued on a finite-dimensional configuration space are the main objects in the theory of classical stochastic processes. These are real vectors with positive entries that sum to one. Geometrically, they form a simplex, a compact convex subset of a real Euclidean vector space [BeZy2006, § 2].
If the stochastic process is a differentiable Markov chain on a $d$-dimensional Euclidean configuration space, then the family of probability vectors $\mathsf{p}(t)$, that describes the state of the process as $t$ increases, satisfies a first-order system of differential equations of the form
$\dot{\mathsf{p}}(t)=\mathsf{K}(t)\,\mathsf{p}(t)\,.$ (1)
We refer to the family of $d\,\times\,d$-matrices $\mathsf{K}(t)$ as the (time-dependent) generator of the dynamics. We require $\mathsf{K}(t)$ to depend on the time variable $t$ in a sufficiently regular manner to satisfy the conditions of global existence of solutions [AdrL1995, § 1]. In addition, we require $\mathsf{K}(t)$ to be column-diagonally dominated with positive off-diagonal entries, that is,
$\mathsf{K}_{i,j}(t)\begin{cases}\,\geq\,0&\forall\,i\neq j\\[8.5359pt] =-\sum_{k\neq i}^{d}\mathsf{K}_{k,i}(t)&\forall\,i=j\end{cases}$ (2)
Under these conditions, (1) specifies a classical master equation governing the forward in time evolution of probability vectors of the Markov chain. In particular, for each $t\,\geq\,s$ the matrix elements of the time-ordered exponential
$\mathsf{T}(t,s)=\mathcal{T}\exp\left(\int_{s}^{t}\mathrm{d}u\,\mathsf{K}(u)\right)$ (3)
solving (1) with initial condition specified by the identity matrix in $\mathcal{M}_{d}(\mathbb{C})$
$\mathsf{T}(s,s)=\mathsf{1}_{d}$ (4)
specify the transition probabilities to go from any state at the earlier time $s$ to another state at the later time $t$. The theory of differential equations guarantees that the maps (3) collectively enjoy the semigroup property
$\mathsf{T}(t_{3},t_{1})=\mathsf{T}(t_{3},t_{2})\mathsf{T}(t_{2},t_{1})\qquad\forall\,t_{3}\,\geq\,t_{2}\,\geq\,t_{1}$ (5)
which guarantees that the Chapman-Kolmogorov equation, a necessary consequence of the Markov property, is satisfied [PavG2014, § 2]. This means that the transition probability matrix (3) maps the simplex of probability vectors in $\mathbb{R}^{d}$ into itself. We thus say that (3) defines a family of classical positive linear maps. A further direct consequence of the interpretation of (3) as transition probabilities is that the dynamics generated by (1) is contractive with respect to the $1$-norm
$\|\mathsf{T}(t,s)(\mathsf{p}(s)-\tilde{\mathsf{p}}(s))\|_{1}\,\leq\,\left\|\mathsf{p}(s)-\tilde{\mathsf{p}}(s)\right\|_{1}\,\equiv\,\sum_{i=1}^{d}|\mathsf{p}_{i}(s)-\tilde{\mathsf{p}}_{i}(s)|\,.$ (6)
Combining contractivity with the property of mapping a convex compact set into itself paves the way for the application of Perron-Frobenius theorem (see e.g. [SenE2006, § 1]). The upshot is that the spectrum of any of the elements of the semigroup (3) is confined to the unit disk of the complex plane, and it is symmetric with respect to the real axis. Furthermore, there is at least one eigenvalue equal to one. The right eigenoperator associated with this eigenvalue is a state operator after appropriate normalization.
2.2 Quantum master equations
State operators are the main objects in the theory of open quantum systems. In the case of a Hilbert space with $d$ dimensions, they are positive semi-definite (henceforth just called “positive”) $d\,\times\,d$ matrices with unit trace. Geometrically, they also form a compact and convex subspace of the matrix space $\mathcal{M}_{d}(\mathbb{C})$ of $d\,\times\,d$ matrices with complex entries [BeZy2006, § 8.4].
Based only on the requirements of trace and self-adjointness preservation, it is in general possible to derive the matrix valued equation
$\dot{\bm{\rho}}(t)=\mathfrak{L}(t)[\bm{\rho}(t)]\,.$ (7)
The action of the generator $\mathfrak{L}(t)$ on any $\operatorname{O}\in\mathcal{M}_{d}(\mathbb{C})$ is always amenable to the following canonical form (see, e.g., [HaCrLiAn2014])
$\begin{split}&\mathfrak{L}(t)[\operatorname{O}]=-\imath\,\left[\operatorname{H}(t),\operatorname{O}\right]+\frac{1}{2}\sum_{\ell=1}^{d^{2}-1}\mathscr{c}_{\ell}(t)\Big(\left[\operatorname{L}_{\ell}(t)\,,\operatorname{O}\operatorname{L}_{\ell}^{\dagger}(t)\right]+\left[\operatorname{L}_{\ell}(t)\operatorname{O}\,,\operatorname{L}_{\ell}^{\dagger}(t)\right]\Big)\\ &\begin{cases}\operatorname{Tr}\left(\operatorname{L}_{\ell}^{\dagger}(t)\operatorname{L}_{k}(t)\right)=\delta_{\ell,k}\\ \operatorname{Tr}\left(\operatorname{L}_{\ell}(t)\right)=0\\ \operatorname{H}(t)=\operatorname{H}^{\dagger}(t)\end{cases}\hskip 14.22636pt\forall\,t\,\geq\,0\hskip 14.22636pt\&\hskip 14.22636pt\forall\,\ell,k=1,\dots,d^{2}-1\end{split}$ (8)
where $\operatorname{H}(t)$ is a Hamilton operator, the $\mathscr{c}_{\ell}(t)$’s are scalar couplings, and the $\operatorname{L}_{\ell}(t)$’s are commonly referred to as decoherence or jump operators. As in the classical case, here we always assume that any explicit time dependence is sufficiently regular, e.g. continuous and uniformly bounded, to satisfy the conditions for global existence of solutions of (7) for any positive time $t\,\geq\,0$.
We refer to (7), always implying the generator form (8), as “quantum master equation”. The main problem with (7) is that it does not ensure the preservation of positivity without further assumptions on the scalar couplings $\mathscr{c}_{\ell}(t)$.
A cornerstone result of the theory of open quantum systems is the proof, see e.g. [RiHu2012, Theorem 4.2.1], that if and only if
$\mathscr{c}_{\ell}(t)\,\geq\,0\qquad\forall\,\ell=1,\dots,d^{2}-1$ (9)
then solutions of (7) can always be couched in the readily positivity-preserving operator sum form (often referred to as the Kraus form [KraK1983])
$\Phi(t,s)[\bm{\rho}(s)]=\sum_{\ell=0}^{d^{2}-1}\operatorname{V}_{\ell}(t,s)\bm{\rho}(s)\operatorname{V}_{\ell}^{\dagger}(t,s)\qquad\forall\,t\,\geq\,s$ (10)
The collection of matrices $\operatorname{V}_{\ell}(t,s)\in\mathcal{M}_{d}(\mathbb{C})$ is such that (10) reduces to the identity map for $t=s$ and
$\sum_{\ell=0}^{d^{2}-1}\operatorname{V}_{\ell}^{\dagger}(t,s)\operatorname{V}_{\ell}(t,s)=\operatorname{1}_{d}\qquad\forall\,t\,\geq\,s$ (11)
to ensure trace preservation.
The proof of (9)-(10) was first obtained in groundbreaking papers by Lindblad [LinG1976] and Gorini-Kossakowski-Sudarshan [GoKoSu1976] under the hypothesis of a time-autonomous generator. This means the case where in (8) the Hamilton operator $\operatorname{H}$, the scalar couplings $\mathscr{c}_{\ell}$’s, and decoherence operators $\operatorname{L}_{\ell}$’s are all time independent. In the time-autonomous case (9) straightforwardly reduces to the requirement that the $\mathscr{c}_{\ell}$’s are positive real numbers:
$\dot{\bm{\rho}}(t)=\mathfrak{L}[\bm{\rho}(t)]:=-\imath\,\left[\operatorname{H},\bm{\rho}(t)\right]+\frac{1}{2}\sum_{\ell=1}^{d^{2}-1}\mathscr{c}_{\ell}\Big(\left[\operatorname{L}_{\ell}\,,\bm{\rho}(t)\operatorname{L}_{\ell}^{\dagger}\right]+\left[\operatorname{L}_{\ell}\bm{\rho}(t)\,,\operatorname{L}_{\ell}^{\dagger}\right]\Big),\ \ c_{\ell}\geq 0.$ (12)
As a consequence, solutions of the master equation only depend upon the time elapsed from the moment an initial state operator $\bm{\rho}$ is assigned:
$\bm{\rho}(t)=\Phi(t,0)[\bm{\rho}],$ (13)
and
$\Phi(t,s)=\Phi(t-s,0)=e^{(t-s)\mathfrak{L}}.$ (14)
Correspondingly, we can regard the collection of linear maps $\Phi$ and matrices $\operatorname{V}_{\ell}$’s as forming a one parameter family, also called homogeneous, with respect to the time variable. The time-autonomous case is of particular interest for applications because it describes quantum master equations obtained in the van-Hove scaling limit [DavE1974].
It is worth recalling that we can always couch the matrix equation (7) in the form of a vector equation using the vectorization isomorphism which relates the outer and tensor products [BeZy2006, § 8.2]. Accordingly, a solution of (7) is in a one-to-one correspondence with a family of vectors in $\mathbb{C}^{d^{2}}=\mathbb{C}^{d}\otimes\mathbb{C}^{d}$
$\bm{\rho}(t)=\sum_{i,j=1}^{d}\mathbf{e}_{i}^{\scriptscriptstyle{(d)}}\mathbf{e}_{j}^{\scriptscriptstyle{(d)}}{}^{\dagger}\rho_{i,j}(t)\quad\Longleftrightarrow\quad\mathscr{r}(t)=\sum_{i,j=1}^{d}\mathbf{e}_{i}^{\scriptscriptstyle{(d)}}\otimes\mathbf{e}_{j}^{\scriptscriptstyle{(d)}}\rho_{i,j}(t)$ (15)
where $\left\{\mathbf{e}_{i}^{\scriptscriptstyle{(d)}}\right\}_{i=1}^{d}$ is the canonical (or computational) basis of $\mathbb{C}^{d}$. In other words, (15) says that we can construct a vector from a matrix by stacking rows one above the other. Upon applying this prescription and using the overline symbol $\overline{(\cdot)}\,$ to denote entry-wise complex conjugation, the evolution map (10) takes the familiar form in linear dynamical systems theory (see, e.g., [CeCeVu2009]) of a matrix acting on a vector:
$\mathscr{F}(t,s)\mathscr{r}(s)=\left(\sum_{\ell=0}^{d^{2}-1}\operatorname{V}_{\ell}(t,s)\otimes\overline{\operatorname{V}_{\ell}}(t,s)\right)\mathscr{r}(s)\qquad\forall\,t\,\geq\,s$ (16)
For any fixed values of $t$ and $s$, $\mathscr{F}(t,s)$ is a $d^{2}\,\times\,d^{2}$ complex square matrix:
$\mathscr{F}(t,s)\in\mathcal{M}_{d^{2}}(\mathbb{C})=\mathcal{M}_{d}(\mathbb{C})\otimes\mathcal{M}_{d}(\mathbb{C})$ (17)
As we vary $t$ and $s$, the $\mathscr{F}(t,s)$’s specify the elements of the semigroup solving the “reshaped” master equation [BeZy2006, § 10.2]
$\begin{split}&\partial_{t}\mathscr{F}(t,s)=\mathscr{L}(t)\mathscr{F}(t,s)\\ &\mathscr{F}(s,s)=\mathsf{1}_{d^{2}}\end{split}\qquad\forall\,t\,\geq\,s$ (18)
Here, $\mathsf{1}_{d^{2}}$ is the identity matrix in $\mathcal{M}_{d^{2}}(\mathbb{C})$ and
$\mathscr{L}(t)$ $=-\imath\,\Big(\operatorname{H}(t)\otimes\mathsf{1}_{d}-\mathsf{1}_{d}\otimes\operatorname{H}^{\top}(t)\Big)$
$+\sum_{\ell=1}^{d^{2}-1}\mathscr{c}_{\ell}(t)\left(\operatorname{L}_{\ell}(t)\otimes\overline{\operatorname{L}_{\ell}}(t)-\frac{\operatorname{L}_{\ell}^{\dagger}(t)\operatorname{L}_{\ell}(t)\otimes\mathsf{1}_{d}+\mathsf{1}_{d}\otimes\operatorname{L}_{\ell}^{\top}(t)\overline{\operatorname{L}_{\ell}}(t)}{2}\right)$ (19)
We use the superscript $\top$ to denote matrix transposition. The one-to-one correspondence between (18) and (9) implies that the operator sum (10) also enjoys the semigroup property:
$\Phi(t_{3},t_{2})\big[\Phi(t_{2},t_{1})[\operatorname{O}]\big]=\Phi(t_{3},t_{1})[\operatorname{O}]\qquad\forall\,t_{3}\,\geq\,t_{2}\,\geq\,t_{1}\quad\&\quad\forall\,\operatorname{O}\in\mathcal{M}_{d}(\mathbb{C})$ (20)
The semigroup generated by the quantum master equation (7) subject to the conditions (9) always preserves the state operators. All these facts combined are reminiscent of the properties of semigroups generated by the classical master equation (1). For this reason quantum master equations satisfying (9) are often called “Markovian”. The set of positive maps in quantum theory is, however, much more involved than in the classical case. The positive operator sum form (10) is only a sufficient condition for positivity preservation. In particular, (10) is equivalent to adding to positivity preservation the requirement [StiW1955, ChoM1975] that the map on $\mathcal{M}_{d^{2}}(\mathbb{C})$ obtained by the tensor product of (10) with the identity on $\mathcal{M}_{d}(\mathbb{C})$ is also positive:
$\tilde{\Phi}^{(d)}(t,s)[\tilde{\operatorname{O}}]:=(\mathrm{Id}_{d}\otimes\Phi(t,s))[\tilde{\operatorname{O}}]\,\geq\,0\qquad\forall\,t\,\geq\,s$ (21)
for all positive matrices $\tilde{\operatorname{O}}$ in $\mathcal{M}_{d^{2}}(\mathbb{C})$. The condition (21) is called $d$- or complete positivity. As a consequence, the master equation (7) subject to the Lindblad-Gorini-Kossakowski-Sudarshan conditions (9) may be more descriptively called completely positive rather than Markovian with reference to the property of the solution semigroup [BaNaTh2008].
Inspecting (21) immediately suggests the existence of a hierarchy of $k$-positive maps $\tilde{\Phi}^{(k)}(t,s)$, $k=1,\dots,d$ characterized by the conditions
$\tilde{\Phi}^{(k)}(t,s)[\tilde{\operatorname{O}}]:=(\mathrm{Id}_{k}\otimes\Phi(t,s))[\tilde{\operatorname{O}}]\,\geq\,0\qquad\forall\,t\,\geq\,s$ (22)
where now $\mathrm{Id}_{k}$ is the identity on $\mathcal{M}_{k}(\mathbb{C})$ and the inequality holds for all positive matrices $\tilde{\operatorname{O}}$ in $\mathcal{M}_{kd}(\mathbb{C})=\mathcal{M}_{k}(\mathbb{C})\otimes\mathcal{M}_{d}(\mathbb{C})$. If we then denote by $\mathcal{P}_{k}$ the set of $k$-positive maps, it is rather straightforward to prove the chain of inclusions [ChrD2022]
$\mathcal{P}\,\equiv\,\mathcal{P}_{1}\supset\mathcal{P}_{2}\supset\dots\supset\mathcal{P}_{d}\,\equiv\,\mathcal{CP}$ (23)
We thus see that the class of $d$-positive maps, i.e., of completely positive maps ($\mathcal{CP}$), is the smallest set in the above chain of inclusions. Unfortunately, for $k$-positive maps, there is no general counterpart of the operator-sum representation (10) that characterizes completely positive maps. Analogously, at the differential level, the conditions (9) have no general counterpart to specify generators of $k$-positive semigroups [WoCi2008, ChrD2022]. Even if it can be argued that completely positive maps are the only ones adapted to describe the evolution of open quantum systems (see, however, the debate [PecP1994, AliR1995, PecP1995] and the more recent [DoLi2016]), completely positive dynamical maps can be solutions of a quantum master equation (7) whose generator does not satisfy the complete positivity conditions (9). The upshot is that quantum master equations that do not satisfy (9) must necessarily be considered. We refer to [SmVa2010] for an example with a clear physical interpretation. In addition, the study of $k$-positive maps for each $k$ is relevant to criteria for discriminating between entangled and separable states (see [ChKo2009] and the references therein). To summarize:
  1. 1.
    (8) equipped with regularity conditions on the time dependence is the generator of a semigroup of trace and self-adjointness preserving linear maps or, equivalently, of infinitely divisible trace and self-adjointness preserving linear maps;
  2. 2.
    we call the semigroup trace preserving and positive if all its elements are linear maps that send the convex set of state operators into itself;
  3. 3.
    the properties under tensor product composition with the identity map filter the set of positive linear maps into the flag of nested subsets (23). Unfortunately, in general the resulting classification does not lead to an explicit representation;
  4. 4.
    however, when (9) holds true the semigroup solution of a quantum master equation only consists of completely positive maps enjoying the property (21). This property is equivalent [StiW1955, ChoM1975] to the Kraus positive operator sum (10);
  5. 5.
    in the time-autonomous case the semigroup enjoys the time translational invariance property (14): it thus reduces to a one parameter family of linear maps.
We conclude the section by recalling that a trace and self-adjointness preserving linear map $\Phi$ is positive if and only if it is a contraction [KosA1972]
$\left\|\Phi(\operatorname{O})\right\|_{1}\,\leq\,\left\|\operatorname{O}\right\|_{1}\qquad\forall\,\operatorname{O}=\operatorname{O}^{\dagger}\in\mathcal{M}_{d}(\mathbb{C})$ (24)
with respect to the trace norm (also known as Schatten $1$-norm)
$\left\|\operatorname{O}\right\|_{1}:=\operatorname{Tr}\left(\sqrt{\operatorname{O}^{\dagger}\operatorname{O}}\right)$ (25)
The trace norm for self-adjoint operators coincides with the sum of the absolute values of all eigenvalues. Thus, positive linear maps are contractive and transform a convex compact set into itself. Again, these facts allow us to demonstrate that positive maps satisfy the hypotheses of the Perron-Frobenius theorem. The conclusion is that all the eigenvalues belong to the unit disk of the complex plane and are symmetric with respect to the real axis. Furthermore, there is at least one eigenvalue equal to one. The right eigenoperator associated with this eigenvalue is a state operator after appropriate normalization.
3 Embedded classical master equation
As (7) is self-adjointness preserving, solutions that evolve from self-adjoint initial conditions at $t=0$
$\bm{\rho}(0)=\sum_{i=1}^{d}p_{i}\,\bm{\varphi}_{i}\bm{\varphi}_{i}^{\dagger}$ (26)
at any later instant of time are amenable to the diagonal form
$\bm{\rho}(t)=\sum_{i=1}^{d}\wp_{i}(t)\,\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)$ (27)
where both the $\bm{\varphi}_{i}$’s and the $\bm{\phi}_{i}(t)$’s constitute a complete orthonormal basis of $\mathbb{C}^{d}$ and the $p_{i}$’s and $\wp_{i}$’s respectively denote the eigenvalues of $\bm{\rho}(0)$, $\bm{\rho}(t)$.
As orthonormal bases are connected by unitary maps, there exists a family of unitaries such that
$\bm{\phi}_{i}(t)=\mathsf{U}(t)\bm{\varphi}_{i}$ (28)
If we insert (28) into (7) and denote by $\bm{p}$, $\bm{\wp}(t)$ the vectors whose entries are the eigenvalues of (26), (27) respectively, a straightforward computation [TeMa1992, WiTo1999] yields
$\begin{split}&\bm{\dot{\wp}}(t)=\mathsf{W}(t)\,\bm{\wp}(t)\\ &\bm{\wp}(0)=\bm{p}\,.\end{split}$ (29)
At any time $t$ the $d\times d$ matrices $\mathsf{W}(t)$ have the same structure as the generator of the classical master equation (2): now they are column-diagonally dominated with elements
$\mathsf{W}_{i,j}(t)=\left\langle\,\bm{\phi}_{i}(t)\,,\mathfrak{L}(t)[\bm{\phi}_{j}(t)\bm{\phi}_{j}^{\dagger}(t)]\,\bm{\phi}_{i}(t)\,\right\rangle=\begin{cases}\mathsf{R}_{i,j}(t)&\forall\,i\neq j\\[8.5359pt] -\sum_{k\neq j}^{d}\mathsf{R}_{k,j}(t)&\forall\,i=j\end{cases}$ (30)
where $\left\langle\,\cdot\,$$\cdot\,\right\rangle$ is the standard inner product in $\mathbb{C}^{d}$ and the off-diagonal elements are the sum over the corresponding matrix elements of the decoherence operators weighted by the canonical couplings:
$\mathsf{R}_{i,j}(t)=\sum_{\ell=1}^{d^{2}-1}\mathscr{c}_{\ell}(t)\,\left|\left\langle\,\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,,\operatorname{L}_{\ell}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\,\right\rangle\right|^{2}$ (31)
Before continuing, some remarks are in order.
  • In general, the “spectral flow” equations (28), (29) do not provide a constructive method for solving (7) but are only a tool to analyze the kinematic properties of individual solutions. More precisely, the matrix elements of (31) depend on the unitary evolution operator in (28). The only general method to identify the latter requires first solving (7) for a given self-adjoint initial condition (26). This also means that the families of matrices $\mathsf{U}(t)$ and $\mathsf{W}(t)$ depend on the choice of the initial conditions. In other words, the spectral flow equations play the same role for the kinematic analysis of quantum master equations as the current velocity formalism plays for the kinematic analysis of the Fokker-Planck equation [NelE2001, § 13].
  • The derivation of (29) requires almost-everywhere differentiability of eigenvalues and eigenvectors of families of self-adjoint operators satisfying (7). Mathematically, any solution $\mathscr{x}(t)$ of a linear differential equation with respect to a real variable $t$ can be analytically continued to a unique holomorphic function of the independent variable:
    $\overline{\mathscr{x}(z)}=\mathscr{x}(\overline{z})\qquad\forall\,z\in\mathbb{C}\quad\text{ where }\operatorname{Re}z=t\,.$ (32)
    As a consequence, we can analytically continue self-adjoint solutions of (7) to matrix valued holomorphic functions
    $\bm{\rho}^{\dagger}(z)=\bm{\rho}(\bar{z})$ (33)
    in a neighborhood of the real $t$-axis. This fact allows us to invoke Theorem 6.1 in [KatT1982, § 2.6.1], which guarantees that the eigenvalues and spectral projectors of a family of operators satisfying (33) are holomorphic functions and, as such, are continuous and differentiable. In addition [KatT1982, § 2.6.1], eigenvectors can also be chosen to be holomorphic functions, thus ensuring the validity of the construction used in the proof.
3.1 Kossakowski conditions
The representation (27) yields the identity
$\left\|\bm{\rho}(t)\right\|_{1}=\sum_{i=1}^{d}|\wp_{i}(t)|=\left\|\bm{\wp}(t)\right\|_{1}$ (34)
holding for any solution of (7) with self-adjoint initial data. This means that (7) is contractive if and only if all classical master equations embedded in particular solutions are contractive.
It is readily clear that if the complete positivity conditions (9) hold true, then the off-diagonal entries (31) are positive functions of time. Consequently, (29) is a classical master equation and, as such, describes contractive dynamics. Conversely, if the entries (31) are always positive for any solution of (7) corresponding to self-adjoint initial data, then the semigroup generated by (7) is necessarily contractive and, therefore, positive [ChrD2022]. In fact, requiring that the (31) are always positive is equivalent to the Kossakowski conditions [KosA1972, GoKoSu1976] (see also [RiHu2012, § 4.3]) asserting that a linear map of the form (8) is the generator of a trace-preserving contraction semigroup if and only if its matrix entries evaluated on an arbitrary orthonormal frame specify a classical generator. More explicitly, if we denote by $\big\{\operatorname{P}_{i}\big\}_{i=1}^{d}$ a complete set of orthonormal projectors on $\mathbb{C}^{d}$, that is,
$\begin{split}&\sum_{i=1}^{d}\operatorname{P}_{i}=\operatorname{1}_{d}\\ &\operatorname{P}_{i}\operatorname{P}_{j}=\delta_{i,j}\operatorname{P}_{i}\end{split}$ $\qquad\forall\,i,j=1,\dots,d$ (35)
then the Kossakowski conditions read
$\begin{cases}\operatorname{Tr}\left(\operatorname{P}_{i}\mathfrak{L}(t)[\operatorname{P}_{j}]\right)\,\geq\,0&i\neq j\\[8.5359pt] \operatorname{Tr}\left(\operatorname{P}_{i}\mathfrak{L}(t)[\operatorname{P}_{i}]\right)=-\sum_{k\neq i=1}^{d}\operatorname{Tr}\left(\operatorname{P}_{k}\mathfrak{L}(t)[\operatorname{P}_{i}]\right)\,\leq\,0&i=j\end{cases}$ (36)
To summarize: the identity (34) allows us to explicitly relate the contractivity of the semigroup generated by (7) to that of solutions of the classical master equation (29) and, therefore, to derive Kossakowski’s conditions.
4 Conditions on the embedded classical master equation stemming from the positivity class of the semigroup
We now turn to derive some properties of (29) that follow if we assume that the semigroup, in addition to being trace preserving and positive, enjoys further properties. To this end, we recall that the Hilbert-Schmidt inner product of two matrices is defined by
$\langle\operatorname{A},\operatorname{B}\rangle:=\operatorname{Tr}\!\left(\operatorname{A}^{\dagger}\operatorname{B}\right)\qquad\forall\,\operatorname{A},\operatorname{B}\in\mathcal{M}_{d}(\mathbb{C})$ (37)
Accordingly, the trace of $\mathfrak{L}(t)$ is
$\operatorname{Tr}\mathfrak{L}(t):=\sum_{\mu=1}^{d^{2}}\langle\operatorname{A}_{\mu}\mathfrak{L}(t)[\operatorname{A}_{\mu}]\rangle$ (38)
where $\{\operatorname{A}_{\mu}\}_{\mu=1}^{d^{2}}$ is a complete orthonormal basis of $\mathcal{M}_{d}(\mathbb{C})$. In particular, for any complete orthonormal basis $\{\bm{f}_{i}\}_{i=1}^{d}$ of $\mathbb{C}^{d}$, the collection of outer products $\{\bm{f}_{i}\bm{f}_{j}^{\dagger}\}_{i,j=1}^{d}$ constitutes a complete orthonormal basis of $\mathcal{M}_{d}(\mathbb{C})$. Hence, we obtain
$\operatorname{Tr}\mathfrak{L}(t)$ $=\sum_{i,j=1}^{d}\operatorname{Tr}\!\left((\bm{f}_{i}\bm{f}_{j}^{\dagger})^{\dagger}\mathfrak{L}(t)[\bm{f}_{i}\bm{f}_{j}^{\dagger}]\right)$
$=\sum_{i,j=1}^{d}\left\langle\bm{f}_{i},\mathfrak{L}(t)[\bm{f}_{i}\bm{f}_{j}^{\dagger}]\,\bm{f}_{j}\right\rangle$ (39)
It is straightforward to verify that (39) is also equal to the trace of the reshaped generator $\mathscr{L}(t)$ defined by (19):
$\operatorname{Tr}\mathfrak{L}(t)=\operatorname{Tr}\mathscr{L}(t)=-d\,\sum_{\ell=1}^{d^{2}-1}\mathscr{c}_{\ell}(t)\,\leq\,0$ (40)
This result implies contractivity of the volume element for any $t\,\geq\,s$:
$\det\mathscr{F}(t,s)=\exp\left(\int_{s}^{t}\mathrm{d}u\,\operatorname{Tr}\mathscr{L}(u)\right)$ (41)
where $\mathscr{F}(t,s)$ is the solution of (18).
4.1 Condition on the embedded classical master equation imposed by 2-positivity
A linear map
$\bm{X}\colon\mathcal{M}_{d}(\mathbb{C})\mapsto\mathcal{M}_{d}(\mathbb{C})$ (42)
is called $2$-positive if for any positive operator $\operatorname{O}$ belonging to $\mathcal{M}_{2d}(\mathbb{C})$
$\operatorname{O}\,\geq\,0\quad\Longrightarrow\quad(\mathrm{Id}_{2}\otimes\bm{X})(\operatorname{O})\,\geq\,0\,.$ (43)
Now, let us assume that (7) generates a $2$-positive semigroup. Then we claim that the classical master equation constructed out of any of the particular solutions is such that
$0\,\geq\,\operatorname{Tr}\mathsf{W}(t)\,\geq\,\frac{1}{d}\operatorname{Tr}\mathfrak{L}(t)$ (44)
An equivalent version of this claim was first proven in [vEnChKiMG2025].
4.1.1 Proof of the claim
The action of a semigroup element close to the identity on a state operator $\bm{\rho}$ admits the explicit representation:
$\Phi(t+\varepsilon,t)[\bm{\rho}]=\bm{\rho}+\varepsilon\,\mathfrak{L}(t)[\bm{\rho}]+O(\varepsilon^{2})$ (45)
Let us denote by $\mathbf{e}_{i}^{\scriptscriptstyle{(2)}}$, $i=1,2$ the elements of the canonical (or computational) basis of $\mathbb{C}^{2}$. We can then construct a vector in $\mathbb{C}^{2d}$ of the form
$\bm{u}=\frac{\mathbf{e}_{1}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}_{i}(t)+\mathbf{e}_{2}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}_{j}(t)}{\sqrt{2}}$ (46)
where $\left\{\bm{\phi}_{i}(t)\right\}_{i=1}^{d}$ is a complete orthonormal basis of $\mathbb{C}^{d}$ determined by the eigenvector flow equation (28). The projector along $\bm{u}$ is clearly a positive matrix
$\operatorname{P}_{u}=\bm{u}\bm{u}^{\dagger}=\frac{1}{2}\begin{bmatrix}\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)&\bm{\phi}_{i}(t)\bm{\phi}_{j}^{\dagger}(t)\\ \bm{\phi}_{j}(t)\bm{\phi}_{i}^{\dagger}(t)&\bm{\phi}_{j}(t)\bm{\phi}_{j}^{\dagger}(t)\end{bmatrix}$. (47)
Then we find
$(\mathrm{Id}_{2}\otimes\Phi(t+\varepsilon,t))[\operatorname{P}_{u}]$ $=\frac{1}{2}\begin{bmatrix}\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)&\bm{\phi}_{i}(t)\bm{\phi}_{j}^{\dagger}(t)\\ \bm{\phi}_{j}(t)\bm{\phi}_{i}^{\dagger}(t)&\bm{\phi}_{j}(t)\bm{\phi}_{j}^{\dagger}(t)\end{bmatrix}$
$+\frac{\varepsilon}{2}\begin{bmatrix}\mathfrak{L}(t)\big[\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)\big]&\mathfrak{L}(t)\big[\bm{\phi}_{i}(t)\bm{\phi}_{j}^{\dagger}(t)\big]\\ \mathfrak{L}(t)\big[\bm{\phi}_{j}(t)\bm{\phi}_{i}^{\dagger}(t)\big]&\mathfrak{L}(t)\big[\bm{\phi}_{j}(t)\bm{\phi}_{j}^{\dagger}(t)\big]\end{bmatrix}+O(\varepsilon^{2})\,.$ (48)
As the semigroup is $2$-positive by assumption, we must have
$\left\langle\,\bm{v}\,,(\mathrm{Id}_{2}\otimes\Phi(t+\varepsilon,t))[\operatorname{P}_{u}]\,\bm{v}\,\right\rangle\,\geq\,0$ $\forall\,\bm{v}\in\,\mathbb{C}^{2d}$ (49)
In particular, we can choose
$\bm{v}=\frac{\mathbf{e}_{1}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}_{i}(t)-\mathbf{e}_{2}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}_{j}(t)}{\sqrt{2}}=\frac{1}{\sqrt{2}}\begin{bmatrix}\bm{\phi}_{i}(t)\\ -\bm{\phi}_{j}(t)\end{bmatrix}$ (50)
so that
$\left\langle\,\bm{u}\,,\bm{v}\,\right\rangle=0\,.$ (51)
In such a case, after a straightforward calculation, we obtain
$0\,\leq\,\lim_{\varepsilon\downarrow 0}\frac{\left\langle\,\bm{v}\,,(\mathrm{Id}_{2}\otimes\Phi(t+\varepsilon,t))[\operatorname{P}_{u}]\bm{v}\,\right\rangle}{\varepsilon}$
$\hskip 14.22636pt=\frac{\left\langle\,\bm{\phi}_{i}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)\big]\bm{\phi}_{i}(t)\,\right\rangle+\left\langle\,\bm{\phi}_{j}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)\big]\bm{\phi}_{j}(t)\,\right\rangle}{2}$
$\hskip 14.22636pt-\frac{\left\langle\,\bm{\phi}_{i}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}_{i}(t)\bm{\phi}_{j}^{\dagger}(t)\big]\bm{\phi}_{j}(t)\,\right\rangle+\left\langle\,\bm{\phi}_{j}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}_{j}(t)\bm{\phi}_{i}^{\dagger}(t)\big]\bm{\phi}_{i}(t)\,\right\rangle}{2}$ (52)
Summing over $i,j=1,\dots,d$, the definitions (30) and (40) yield the claim.
4.2 Condition on the embedded classical master equation imposed by the Schwarz-class
It can be shown (see, e.g., the discussion in [ChKiMu2024] and references therein) that the convex set of positive maps on $\mathcal{M}_{d}(\mathbb{C})$ yields a hierarchy of inclusions finer than (23). For the purposes of the present work, we only note that between the convex sets of positive and $2$-positive maps lies the set $\mathcal{S}_{1}$ (also convex) of Schwarz maps:
$\mathcal{P}\,\equiv\,\mathcal{P}_{1}\supset\mathcal{S}_{1}\supset\mathcal{P}_{2}$ (53)
We characterize Schwarz maps by considering their adjoint with respect to the Hilbert-Schmidt inner product (37). For an arbitrary linear map $\bm{X}$, we define its Hilbert-Schmidt adjoint $\bm{X}^{\ddagger}$ by requiring
$\langle\operatorname{A}\,,\bm{X}[\operatorname{B}]\rangle:=\operatorname{Tr}\left(\operatorname{A}^{\dagger}\bm{X}[\operatorname{B}]\right)=\operatorname{Tr}\left(\big(\bm{X}^{\ddagger}[\operatorname{A}]\big)^{\dagger}\operatorname{B}\right)=:\langle\bm{X}^{\ddagger}[\operatorname{A}]\,,\operatorname{B}\rangle\qquad\forall\,\operatorname{A},\operatorname{B}\in\mathcal{M}_{d}(\mathbb{C})$ (54)
It is straightforward to verify that trace-preservation imposes the adjoint to be unital
$\bm{X}^{\ddagger}[\operatorname{1}_{d}]=\operatorname{1}_{d}\,.$ (55)
We say that a positive unital map is Schwarz class if
$\bm{X}^{\ddagger}[\operatorname{A}^{\dagger}\operatorname{A}]\,\geq\,\bm{X}^{\ddagger}[\operatorname{A}^{\dagger}]\bm{X}^{\ddagger}[\operatorname{A}]\qquad\forall\,\operatorname{A}\,\in\,\mathcal{M}_{d}(\mathbb{C})$ (56)
It is worth recalling here that Kadison’s theorem asserts that the same inequality is satisfied by all positive maps if the domain is restricted to normal operators [KadR1952]. Conversely, $2$-positivity of unital maps implies the Schwarz property.
Let us assume that a semigroup is Schwarz class. A straightforward calculation shows that the generator of such a semigroup must satisfy the dissipativity condition:
$\mathfrak{L}^{\ddagger}(t)[\operatorname{A}^{\dagger}\operatorname{A}]\,\geq\,\mathfrak{L}^{\ddagger}(t)[\operatorname{A}^{\dagger}]\operatorname{A}+\operatorname{A}^{\dagger}\mathfrak{L}^{\ddagger}(t)[\operatorname{A}]\qquad\forall\,t\quad\&\quad\forall\,\,\operatorname{A}\,\in\,\mathcal{M}_{d}(\mathbb{C})\,.$ (57)
We refer to [ChKiMu2024] for explicit examples.
Now, let us suppose that (7) generates a semigroup in $\mathcal{S}_{1}$. Then we claim that the classical master equation constructed from any of the particular solutions is such that
$0\,\geq\,\operatorname{Tr}\mathsf{W}(t)\,\geq\,\frac{2}{d+1}\operatorname{Tr}\mathscr{L}(t)\,.$ (58)
An equivalent version of this claim was first proven in [ChvEnKiMG2025].
4.2.1 Proof of the claim
If we evaluate the dissipativity condition (57) on non-self-adjoint projectors constructed with eigenvectors of (27) at time $t$
$\operatorname{P}_{i,j}(t)=\bm{\phi}_{i}(t)\bm{\phi}_{j}^{\dagger}(t)$ (59)
we get
$\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{j,j}(t)]\,\geq\,\operatorname{P}_{i,j}^{\dagger}(t)\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}(t)]+\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}^{\dagger}(t)]\operatorname{P}_{i,j}(t)\,.$ (60)
From the left-hand side of the inequality we obtain:
$\sum_{i=1}^{d}\sum_{j\neq i}^{d}\operatorname{Tr}\left(\operatorname{P}_{j,j}(t)\mathfrak{L}^{\ddagger}(t)\big[\operatorname{P}_{j,j}(t)\big]\right)=\sum_{i,j=1}^{d}(1-\delta_{i,j})\operatorname{Tr}\left(\mathfrak{L}(t)\big[\operatorname{P}_{j,j}(t)\big]\operatorname{P}_{j,j}(t)\right)=(d-1)\,\operatorname{Tr}\mathsf{W}(t)$ (61)
The result must be larger than or equal to the corresponding expression for the right-hand side of (57)
$\sum_{i=1}^{d}\sum_{j\neq i}^{d}\operatorname{Tr}\left(\operatorname{P}_{j,j}(t)\big(\operatorname{P}_{i,j}^{\dagger}(t)\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}(t)]+\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}^{\dagger}(t)]\operatorname{P}_{i,j}(t)\big)\right)$
$=\sum_{i,j=1}^{d}(1-\delta_{i,j})\operatorname{Tr}\left(\operatorname{P}_{j,i}(t)\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}(t)]+\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}^{\dagger}(t)]\operatorname{P}_{i,j}(t)\right)=2\,\left(\operatorname{Tr}\mathscr{L}(t)-\operatorname{Tr}\mathsf{W}(t)\right)$ (62)
The claim follows at once by comparing the two identities above.
5 A distilled view on Lyapunov exponents
We briefly recall the definition and the properties of Lyapunov exponents propaedeutic to the proof of our main result. In doing so, we mainly draw from [AdrL1995] and the more recent [BarL2017].
Let $\mathsf{X}(t)$ be the Cauchy matrix solution of a time-non-autonomous differential system on a $d$-dimensional Euclidean vector space $\mathbb{E}$ with the initial condition given by the identity matrix on $\mathbb{E}$:
$\begin{split}&\dot{\mathsf{X}}(t)=\mathsf{A}(t)\mathsf{X}(t)\\ &\mathsf{X}(0)=\mathsf{1}_{d}\end{split}$ (63)
We assume that $\mathsf{A}(t)$ is a continuous and uniformly bounded square matrix for all $t\,\geq\,0$. Under these hypotheses, if $\big\{\bm{g}_{i}\big\}_{i=1}^{d}$ is any basis of $\mathbb{E}$, then the characteristic exponents
$\chi(\bm{g}_{i})=\limsup_{t\to\infty}\frac{1}{t}\ln\|\mathsf{X}(t)\bm{g}_{i}\|_{a}$ $i=1,\dots,d$ (64)
are well defined quantities. The definition is independent of the choice of the norm $\left\|\cdot\right\|_{a}$ provided the norm is equivalent to the Euclidean norm which corresponds to $a$ equal to $2$. We elaborate some more on norm equivalence in Appendix LABEL:app:norms.
Lyapunov introduced the term “normal basis” for a basis $\big\{\bm{f}_{i}\big\}_{i=1}^{d}$ that minimizes the sum over all the characteristic exponents [OseV1968, BeGaGiSt1980]:
$\sum_{i=1}^{d}\chi(\bm{f}_{i})=\liminf_{t\to\infty}\frac{1}{t}\ln\det\mathsf{X}(t)=\liminf_{t\to\infty}\frac{1}{t}\int_{0}^{t}\mathrm{d}s\operatorname{Tr}\mathsf{A}(s)$ (65)
Correspondingly, the exponents (66) computed on a normal basis are called Lyapunov exponents:
$\lambda_{i}=\chi(\bm{f}_{i})$ $i=1,\dots,d$ (66)
Geometrically, a normal basis filters $\mathbb{E}$ into a flag of $n\,\leq\,d$ nested linear subspaces
$\{0\}:=\mathbb{V}_{0}\subset\mathbb{V}_{1}\subset\dots\subset\mathbb{V}_{n}=\mathbb{E}$ (67)
each of which is spanned by a subset of the elements of the basis. This means that each $\mathbb{V}_{i}$ corresponds to a distinct value of the Lyapunov exponents. Clearly, the flag has $d$ elements if the Lyapunov exponents are non-degenerate. In what follows, we label the exponents from one to $d$ to simplify the notation.
The QR decomposition of $\mathsf{X}(t)$, see e.g. [HoJo2013, Th 2.1.14], gives a more computationally concrete meaning to Lyapunov exponents. Namely, it is always possible to uniquely factorize the solution $\mathsf{X}(t)$ of (63) into a unitary matrix $\mathsf{U}(t)$ and an upper triangular matrix $\mathsf{T}(t)$ having positive diagonal entries:
$\mathsf{X}(t)=\mathsf{U}(t)\,\mathsf{T}(t)$ (68)
Upon denoting by $\mathsf{T}_{i,i}(t)$ the diagonal entries of $\mathsf{T}(t)$, we arrive at
$\lambda_{i}=\limsup_{t\to\infty}\frac{1}{t}\ln\mathsf{T}_{i,i}(t)$ $\forall\,i=1,\dots,d$ (69)
[…]
$\begin{split}&\dot{\mathsf{Y}}(t)=-\mathsf{A}^{\dagger}(t)\mathsf{Y}(t)\\ &\mathsf{Y}(0)=\mathsf{1}_{d}\end{split}$ (70)
Traditionally see e.g. [ByViGrNe1966, § 1.4] or [AdrL1995, § I.1], (70) is referred to as the “adjoint equation” to (63). To avoid confusion with the concept of “adjoint” used in section 4.2, we refer to (70) as the Lyapunov “dual system”. The Cauchy matrix solving (70) is the inverse of the adjoint of that solving (63).
[…]
The important consequence is the preservation of the inner product
$\left\langle\,\mathsf{Y}(t)\bm{w}\,,\mathsf{X}(t)\bm{v}\,\right\rangle=\left\langle\,\bm{w}\,,\bm{v}\,\right\rangle$ $\forall\bm{v}\,,\bm{w}\in\mathbb{E}_{d}$ (72)
Upon applying the Cauchy-Schwarz inequality we get
$\|\mathsf{Y}(t)\bm{w}\|_{2}^{2}\,\|\mathsf{X}(t)\bm{v}\|_{2}^{2}\geq\left|\left\langle\,\bm{w}\,,\bm{v}\,\right\rangle\right|^{2}$ $\forall\bm{v}\,,\bm{w}\in\mathbb{E}_{d}$ (73)
Hence, if we arrange the Lyapunov exponents of the primal dynamics (63) in decreasing order and those of the dual dynamics (70) in increasing order, and we choose $\bm{v}$ and $\bm{w}$ as the corresponding elements of normal bases spanning the two flags of nested subspaces, we generally obtain the inequalities:
$\lambda_{i}+\Gamma_{i}\,\geq\,0$ $i=1,\dots\,d$ (74)
These inequalities raise the question of identifying the conditions under which
$\lambda_{i}=-\Gamma_{i}$ $i=1,\dots,d$ (75)
Cardinal results in the theory of Lyapunov exponents due to Oseledets [OseV1968, OseV2008] and Ragunathan [RagM1979], see also [BeGaGiSt1980], show that answering this question is equivalent to determining the conditions under which we can replace $\limsup$ with $\lim$ in the definition (66), or equivalently (69), of the exponents. The answer identifies the class of Lyapunov “regular” systems, which enjoy the property that the limit
$\lim_{t\to\infty}\Big(\mathsf{X}^{\dagger}(t)\mathsf{X}(t)\Big)^{\frac{1}{2\,t}}=\mathsf{Q}$ (76)
exists and specifies a well-defined, self-adjoint, positive square matrix $\mathsf{Q}$. We refer to [BarL2017, Th 7.2.2] for a proof of the equivalence of the characterization of regularity (76) to the one used in [OseV1968, BeGaGiSt1980] and to [BarL2017, Th 4.3.1] for a proof that Lyapunov regularity is equivalent to (75).
When the limit (76) exists, the Lyapunov exponents coincide with the logarithms of the eigenvalues $\big\{\mathsf{q}_{i}\big\}_{i=1}^{d}$ of $\mathsf{Q}$:
$\lambda_{i}=-\Gamma_{i}=\ln\mathsf{q}_{i}$ $i=1,\dots,d$ (77)
Furthermore, for a regular system, the sum of the Lyapunov exponents exactly specifies the growth rate of a volume form:
$\sum_{i=1}^{d}\lambda_{i}=\lim_{t\to\infty}\frac{1}{t}\int_{0}^{t}\mathrm{d}s\operatorname{Tr}\mathsf{A}(s)$ (78)
The characterization of regularity (76) is the basis of many algorithms used in the physics literature [CeCeVu2009, PiPo2016] to compute Lyapunov exponents. Mathematically, (76) is a stronger requirement than global existence and uniqueness of solutions (63). Fortunately, it is possible to prove sufficient conditions for regularity only based on the properties of the matrix $\mathsf{A}(t)$. Specifically, (63) is Lyapunov regular if $\mathsf{A}(t)$, in addition to be continuous and bounded, enjoys one of the following properties:
  1. C-i
    is time independent
    $\mathsf{A}(t)=\mathsf{A}(0)=\mathsf{A}$ $\forall\,t\,\geq\,0$ (79)
    as we can directly verify by diagonalization see e.g. [AdrL1995, Lemma 3.5.3]);
  2. C-$i$
    is periodic in time with period $T>0$:
    $\mathsf{A}(t+T)=\mathsf{A}(t)$ $\forall t\,\geq\,0$ (80)
    This claim can be verified by means of Floquet theorem see e.g. [AdrL1995, Lemma 3.5.3].
  3. C-iii
    $\mathsf{A}(t)$ is almost reducible to a time independent or periodic matrix. Almost reducible means that for any $\delta\,\geq\,0$ there exists a nonsingular and continuously differentiable matrix $\mathsf{L}_{\delta}(t)$ such that
    $\mathsf{B}(t)+\mathsf{C}(t)=\mathsf{L}_{\delta}^{-1}(t)\mathsf{A}(t)\mathsf{L}_{\delta}(t)-\mathsf{L}_{\delta}^{-1}(t)\mathsf{\dot{L}}_{\delta}(t)$ (81)
    such that $\mathsf{B}(t)$ is either constant or periodic and
    $\|\mathsf{C}(t)\|\,\leq\,\delta$ (82)
    The matrix norm may be any norm induced by a vector norm equivalent to the Euclidean norm. The proof of this last sufficient condition is more laborious [AdrL1995, Th 3.5.2].
The sufficient conditions C-iC-iii encompass time-autonomous, periodic and almost periodic generators of quantum master equations. Generators typically, if not always, encountered in applications involving deterministic quantum master equations satisfy one of the sufficient conditions above.
More generally, Oseledets’ multiplicative ergodic theorem [OseV1968, RagM1979, BeGaGiSt1980, OseV2008] guarantees the existence of the limit (76) when, roughly speaking, the family of Cauchy matrices constitutes a statistically stationary, ergodic sequence of matrices. The theorem is of central importance for non-autonomous linear systems obtained by linearization around a solution $\bm{x}(t)$ of a non-linear differential equation to study its stability versus infinitesimal change of the initial data. In this latter case the hypotheses of Oseledets’ theorem are satisfied e.g. when $\bm{x}(t)$ is ergodic. Physics applications of Oseledets’ theorem to non-linear dynamics are beyond the scopes of the present contribution and are extensively discussed in [CeCeVu2009, PiPo2016]. We expect, however, Oseledets’ theorem to be relevant for studying ensembles of random quantum master equations [DeLaTaChZy2019].
6 Main result
We are now in a position to prove results equivalent to those of [ChvEnKiMG2025, vEnChKiMG2025]. Semigroups generated by (7) are always trace and self-adjointness preserving. In addition
Proposition 6.1
If the elements of the semigroup are all contractive and Lyapunov regular, the Lyapunov exponents satisfy
$0\,=\,\lambda_{0}\,\geq\,\lambda_{1}\,\geq\,\dots\,\geq\,\lambda_{d^{2}-1}$ (83)
while, according to the definition of section 5, the exponents of the dual system are
$\Gamma_{d^{2}-1}=-\lambda_{d^{2}-1}\,\geq\,\dots\,\geq\,\Gamma_{1}=-\lambda_{1}\,\geq\,\Gamma_{0}=-\lambda_{0}=0$ (84)
Furthermore,
  1. 1.
    if the semigroup is also $2$-positive, i.e., (44) holds, then the smallest Lyapunov exponent obeys the bound
    $\lambda_{d^{2}-1}\,\geq\,\frac{1}{d}\sum_{i=1}^{d^{2}-1}\lambda_{i}$ (85)
  2. 2.
    if instead the semigroup is Schwarz class, i.e., (58) holds, then
    $\lambda_{d^{2}-1}\,\geq\,\frac{2}{d+1}\sum_{i=1}^{d^{2}-1}\lambda_{i}$ (86)
  3. 3.
    finally, if the semigroup does not enjoy any additional property, then the lower bound on the smallest Lyapunov exponent is trivial:
    $\lambda_{d^{2}-1}\,\geq\,\sum_{i=1}^{d^{2}-1}\lambda_{i}$ (87)
Before turning to the proof, some observations are in order.
  • Our hypotheses require contractivity for all the elements of the semigroup or, in other words, positive divisibility of the linear maps.
  • The dual exponents $\Gamma_{i}$’s specify the asymptotic decay rates of a contractive semigroup. The inequalities for the asymptotic decay rates read:
    $\Gamma_{d^{2}-1}\leq\varkappa_{d}\sum_{\ell=1}^{d^{2}-1}\Gamma_{\ell}$ (88)
    where the universal prefactor $\varkappa_{d}$ is
    $\varkappa_{d}=\begin{cases}\end{cases}$